If f(x)=x^2/5−5x^−2+2x^0, then what is the value of f(5)?
Answer by Boreal(15235)

f(x)=x^2/5−5x^−2+2x^0,
f(5)=5^2/5-5*(5^(-2)+2(5^0)
=5-(5/25)+2(1)
=7-1/5=34/5 or 6.8.